Your parents will retire in 20 years. They currently have $250,000, and they think they will need $1 million at retirement. What annual interest rate must they earn to reach their goal, assuming they don't save any additional funds? Round your answer to two decimal places.
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

1,000,000=250,000*(1+r/100)^20

(1,000,000/250,000)^(1/20)=(1+r/100)

(1+r/100)=1.0718

r=1.0718-1

=7.18%(Approx).